• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 22:55:59 on August 19, 2008

    Entry number 242221

    keyword=S2m Delay for TDCs Lengthened and ADC Channels Swapped

    With cosmice data, we found that the S2m Right ADC channels were empty.
    This was traced to a problem with the cable plugged into the ADC. I
    found that the support for the VDC ribbon cables restricts the ADC
    cable's path to the first input in Slot 25. I decided to move the S2m
    Right PMT channels from ROC 3, Slot 25 channels 0-15 to channels 48-63.

    I have updated the data base to reflect this change on bbsoftw.

    We also saw that the TDC signals from S2m were at the extreme left (low
    TDC channel). In fact, the right PMT signals were mostly missing. With
    Alexandre and Kalyan's help, we checked the TDC signal timing with
    respect to the TDC common start. The TDC signal time was found to be
    too close to the common start time. We delayed the TDC signal with
    additional ribbon cable of about 40 ns delay for both S2m left and right
    signals. Now the TDC signals are between TDC channels 850 and 1200.
    There are still eight signals that are less than 1000 channels.

    We may need to add more delay. The other option is to adjust the
    retiming signal, but I didn't want to tangle with the retiming without
    consulting Bob first.

    A copy of this log entry has been emailed to: rom@jlab.org,kalyan@jlab.org,camsonne@jlab.org,ycwang@jlab.org,vasulk@jlab.org